I am stuck on correct answer for:
a solution to equation 2x^2 - 7x + 3 = 0 is:

I am not using a specific book. I am just practicing...

Answer by venugopalramana(3286)

I am stuck on correct answer for:
a solution to equation 2x^2 - 7x + 3 = 0 is:
2X^2-6X-X+3=0
2X(X-3)-1(X-3)=0
(X-3)(2X-1)=0
X-3=0...........THAT IS ......X=3
2X-1=0....THAT IS..............X=1/2
I am not using a specific book. I am just practicing...